OpenStudy (anonymous):

[2(-2+h)^{2} +3(-2+h)] -2= ?

OpenStudy (anonymous):

\[[2(-2+h)^{2} +3(-2+h)] -2= ?\] thanks

OpenStudy (paxpolaris):

Do you want to factor:\[2(-2+h)^{2} +3(-2+h) -2\] Assume -2+h=x

OpenStudy (anonymous):

no don't factor, it's only a part of a problem

OpenStudy (anonymous):

just simplify

OpenStudy (anonymous):

fyi, it's the whole bracket minus 2 [2(-2+h)^{2} +3(-2+h)] -2= ?

OpenStudy (paxpolaris):

brackets don't matter ...it's still the same

OpenStudy (anonymous):

oh hehe :)

OpenStudy (paxpolaris):

So just the expanded form is your answer

OpenStudy (paxpolaris):

??

OpenStudy (anonymous):

so what does it come to? my online hw has it coming to -5h+5h^2 how do they get that?

OpenStudy (anonymous):

when i do it, it comes to 3h+2h^2

OpenStudy (paxpolaris):

\[2\left( h^2-4h+4 \right)+(3h-6)-2\]

OpenStudy (paxpolaris):

it can't be 5h^2

OpenStudy (paxpolaris):

or something is wrong with the question you typed above

OpenStudy (anonymous):

my mistake was distributing the square do you want to see the whole problem?

OpenStudy (paxpolaris):

sure

OpenStudy (anonymous):

so f'(a)=lim [f(a+h)-f(a)]/h h-->0

OpenStudy (anonymous):

\[f(x)=2x ^{2}+3x\] a=-2

OpenStudy (anonymous):

\[f'(-2)=\lim h-->0 [(2(-2+h)^{2}+3(-2+h))-2]/h\]

OpenStudy (paxpolaris):

\[\Large f'(a)=\lim_{h \rightarrow 0}{f(a+h)-f(a) \over h}\]... that the definition of the derivative...ok

OpenStudy (anonymous):

yes exactly!

OpenStudy (anonymous):

\[f'(-2)=\lim_{h \rightarrow 0} (2(-2+h)^{?} +3(-2+h))-2\div h\]

OpenStudy (anonymous):

then after that is the answer \[\lim_{h \rightarrow 0} -5h+2h ^{2}\div h\]

OpenStudy (paxpolaris):

that's right

OpenStudy (anonymous):

but how do they get there?

OpenStudy (paxpolaris):

back to your original question: \[2(−2+h)^2+3(−2+h)−2\]\[=(2h^2-8h+8)+(3h-6)-2\]\[=2h^2+(-8h+3h)+(8-6-2)\]

OpenStudy (anonymous):

oh my god yeah i was doing the square wrong

OpenStudy (anonymous):

you are a gentleman, thanks so easy, i'm just not thinking thank you so much

OpenStudy (paxpolaris):

\[so,\ \Large f'(-2)=-5\]